# Multi-modal atmospheric characterization of Pictoris b

### v1

@@ 2.2 Archival data
In addition to GRAVITY data, we used archival spectroscopic and photometric data. Table 2 compiles in more details the characteristics of each observations. Data reduction and post-processing analysis can be found in the referenced paper of each study. The specific analysis of high-resolution data is presented in more details in Sect. 6. We do not include the six new JWST/NIRCam photometric points presented in Kammerer et al. (2024), as the two PSF subtraction methods used in that work produced significant scaling differences in the final flux values. Moreover, we do not expect these additional points to contribute meaningfully to the fit in our current weighted scheme.
@@
| VLT/SPHERE | YJ-band spectroscopy | 0.95-1.35 µm | 50 | 22.5 | yes | (*) |
@@
| Gemini/GPI | YJHK-band spectroscopy | 0.9-2.4 µm | 45 | 10 | yes | (b) |
| VLTI/GRAVITY | K-band spectroscopy | 1.97-2.4 µm | 4 000 | 20 | yes | (**) |
| VLT/SINFONI | K-band spectroscopy | 1.929-2.472 µm | 4 120 | 20 | no | (c) |
| VLT/CRIRES+ | K-band spectroscopy | 2.06-2.47 µm | 100 000 | ¿5 | no | (d, e) |
| M-band spectroscopy | 3.51-5.22 µm | 100 000 | ¿5 | no | ||
| VLTI/MATISSE | LM-band spectroscopy | 2.76-4.99 µm | 387-700 | 35 | yes | (f) |
| JWST/MIRI | M-band spectroscopy | 4.90-6.8 µm | 30-40 | 15 | yes | (g) |
@@ 3 Methods
In this section, we present our fitting strategy, namely forward modeling, as well as our Bayesian code ForMoSA. Subsect. 3.1 and 3.2 deals respectively with the forward models we used and the new additions made to ForMoSA for the purpose of this work.
To extract information from exoplanet spectra, a common approach is to compare them with synthetic spectra generated from models of atmosphere. The comparison heavily relies on Bayesian inversion methods that can be divided into two categories: forward modeling and retrieval. The forward modeling approach (Samland et al., 2017) works ”from the model to the data” by using a pre-computed grid of several thousand of spectra and comparing them to the data using Bayesian inference methods. This approach has the advantage of being ”self-consistent”, meaning that it faithfully maintains the physical interpretation of the observed spectrum, as each synthetic spectrum in the grid is computed with strong physical priors. It is also relatively fast since you do not need to generate a new model at each step, but only need to re-interpolate the grid. The main drawbacks are the dependencies on both the quality and the size of the parameter space proposed by the model grid, as well as the systematic deviations from the models that can arise from a lack of physics or hidden fixed free-parameters in the forward model.
Each model is compared iteratively using Bayesian inversion techniques. The goal is to estimate the posterior probability of the model parameters given the data and prior knowledge on the parameters , using Bayes theorem:
@@
is the probability of the data given a parameter set, also known as the likelihood function. This is the function we want to maximize during the inversion.
is the Bayesian evidence, computed by marginalizing over the parameter space. It can be written with the prior volume.
The Bayesian evidence is especially useful when comparing two competing models and . Unlike the likelihood , it is independent of any single set of parameters as it integrates over all parameters (and is therefore more general in that sense). Concretely, if we name and the Bayesian evidences of the two competing models, we can compute the Bayes factor:
Typically, a Bayes factor above 12 (or above 2.5) can be interpreted as a ”moderate need” for the first model to describe the data instead of the second ( significance, see table 4 of Benneke & Seager, 2013). detection. Recently, Kipping & Benneke (2025) demonstrated that this commonly used correspondence between and can lead to overestimations of the actual detection significance. For completeness, we report both metrics, but we rely primarily on the Bayesian evidence as the main diagnostic tool for model comparison. This approach will be used in Sects. 4.3 and 5 to assess the significance of the 13CO detection. For a more detailed description of the Bayesian inference methods, see Knuth et al. (2015).
For this work, we chose to use the forward modeling approach with our Bayesian inversion tool, ForMoSA. A detailed description of the code can be found in Petrus et al. (2020), Petrus et al. (2021), Petrus et al. (2023), and Petrus et al. (2024). In brief, ForMoSA uses a grid of forward models (see Sect. 3.1) along with spectroscopic and/or photometric data. The tool then proceeds to adapt these models to the data (in terms of wavelength range coverage, resolution, continuum, etc.) to make them comparable. The inversion is performed by maximizing the likelihood function using the nested sampling method (see Skilling, 2004). The grid is interpolated linearly in the N dimension of the parameter space using the python packaged xarray333https://docs.xarray.dev/en/stable/.
@@ 3.1 Forward models
Atmospheric models aim to replicate the emission spectra of exoplanets. To do this, they typically compute a blackbody-like emission spectrum from an interior model, which is then modified by a radiative-convective atmosphere composed of various opacity sources (and sometimes clouds), following a parameterized pressure-temperature profile; all within a simplified 1D framework. These models cover different wavelength ranges and spectral resolutions, as illustrated in Fig. 1. Below is a brief summary of the grids we used, as presented in Petrus et al. (2024):
@@
Exo-REM (Charnay et al., 2018) is our reference model. It includes non-equilibrium chemistry of CO-CH4 and N2-NH3 and takes into account multiple opacity sources: H2-H2 and H2-He collisions induced absorptions, H2O, CH4, CO, CO2, NH3, PH3, TiO, VO and FeH ro-vibrationnal transitions as well as Na and K resonant lines. Clouds formation based on Si, Fe, Na2S, KCl and H2O species is considered but with a simplified cloud micro-physics (i.e. only the fastest chemical processes are taken into account). We used a custom version of the grid to explore the abundance of the isotope 13C resulting in these five free parameters: [1200 K; 2000 K], log(g) [3.5 dex; 4.5 dex], [M/H] [-0.5 dex; 1 dex], C/O [0.40; 0.65] and log(12C/13C) [1.0; 2.6]. Exo-REM is known to produce deviant spectra so we performed a simple cleanup by computing the integrated fluxes across the grids and removing the spectra that did not follow the approximate Stefan–Boltzmann law. Furthermore, we chose to remove all nodes at = 1500 K from the grid as they appeared to have some deviant spectra driving the solutions.
@@
ATMO (Tremblin et al., 2015) are a family of cloudless models. They explore numerous opacity sources : H2-H2 and H2-He collision induced absorptions, NH3, H2O, CH4, CO, TiO and VO ro-vibrationnal transitions as well as Na and K resonant lines. These models include two non-equilibrium chemical reactions: CO-CH4 and N2-NH3 that will drive fingering convection in the atmosphere and change the temperature gradient. This will uniquely emulate the reddening effect of clouds. We also utilise a custom version of this grid exploring 13C. The parameter space explored is: [1600 K; 1900 K], log(g) [3.5 dex; 4.5 dex], [M/H] [-0.6; 0.6], [1.01; 1.05], and log(12C/13C) [1.0; 2.6] ( is the adiabatic index driving the temperature gradient). The grids were set with C/O = 0.55. In Sect. 4, we use an extended grid in and log(g) that explored the following parameters : [800 K; 3000 K], log(g) [2.5 dex; 4.5 dex], [M/H] [-0.6; 0.6] and C/O [0.3; 0.7].
@@
BT-Settl (Allard et al., 2012) model grids explore more complex cloud microphysics than Exo-REM and include non-equilibrium chemistry (for CO, CH4, N2, NH3 and CO2) as well as vertical mixing. To properly take into account the cloud micro-physics, BT-Settl divides the atmosphere into multiple layers and computes the distribution and size of grains by comparing their characteristic times of condensation, coalescence, dispersion and sedimentation in the atmosphere. These clouds are Fe and Si based. We used a custom version of this grid exploring the C/O, giving the following parameter space: [1400 K; 2200 K], log(g) [2.5; 5.0] and C/O [0.2754; 1.096]. The metallicity is solar () for this grid.
@@
Sonora (Morley et al., 2024) models include vertical mixing, multiple opacity sources and parameterized clouds. We used the custom Diamondback sub-model grid that contains, in total, 31 molecules, atoms and ions as opacity sources (hydrocarbons, CO, CO2, OCS, Li, K, Na, LiH, LiF), as well as multiple collision induced absorptions (H2-H2, H2-He, H2-N2, H2-CH4). Clouds are parameterized using Ackerman & Marley (2001) approach, meaning that their opacities are computed for Mie diffusion on MgSiO3, Mg2SiO4 and Al2O3. In the end, the model grid effectively explores the following free parameters: [900 K; 2400 K], log(g) [3.5 dex; 5.5 dex], [M/H] [-0.5; 0.5] and fsed [1; 8] (fsed is the sedimentation efficiency parameter). The grids were set with C/O=0.55.
@@
Atmospheric retrievals and forward modeling approaches both deal with a number of issues when combining different observations into a single Bayesian framework. A non exhaustive list of theses limitations include : data that spectraly overlap (i.e. GPI and GRAVITY); data that have different noise estimates (i.e. spectral covariances or not); data that are not calibrated in flux or have inaccurate flux calibration; data that are taken at different epochs, and spectra for which the spectral continuum can not be trusted or has been removed at previous analysis steps (Doppler spectroscopy, etc). Inspired by the work of Hayoz et al. (2023), we incorporated a Multi-modal Option for Simultaneous Analysis and Improved Constrains (MOSAIC) in ForMoSA aim to reduce theses limitation by adapting and inverting the different datasets in parallel.
The general idea is summarized on Fig. 2. Each dataset () is first imported and adapted alongside a grid of synthetic spectra () taken from a single atmospheric model (). This means that the continuum subtraction and/or wavelength windows can be applied separately for each / pair and then stored for the inversion. Secondly, each is compared with the interpolated using separate likelihood functions during the inversion. The grids are modified by both interpolating them with the original model parameters and then adding transformations with extra-grid parameters that can vary from dataset to dataset (i.e. different scaling factors ck, radial velocities rv, projected rotational velocities v.sin(i), …). This way, we are able to control the transformations we want to apply on a given pair /. The stored sub-grids are usually much smaller than the original grids. Therefore, this step is relatively fast.
New likelihood functions are now available to account for differences in the data reduction scheme and, in particular, noise processing. Given data (spectrum or photometry) , their associated error bars or covariance matrix , and the model vector , the classical distribution is defined as when the observational data points are independent. We implemented in addition when the data covariances are provided as input of the code. We also allowed for various proposed mappings (Zucker, 2003; Brogi & Line, 2019; Ruffio et al., 2019) of cross-correlation signals to likelihood (CCF-to-log) which can be used when Doppler spectroscopy (e.g., Keck/KPIC, VLT/HiRISE, etc) is performed solely or in addition to regular spectrophotometric observations (e.g., JWST, VLT/SPHERE, Gemini/GPI, etc).
We compute the joint likelihood function summing over the likelihoods corresponding to each dataset taken as input. This means that this upgraded version of ForMoSA does not account for correlations either between datasets from the same instrument taken at different epochs or between datasets from different instruments. As is common practice in the field, each dataset is treated as statistically independent.
@@ 4 Inversion of GRAVITY spectra
For this inversion, the likelihood function was computed using the generalized distribution (see Sect. 3.2) taking as input the covariance matrixes provided by the GRAVITY pipeline. The inversion was performed across the entire spectral coverage of GRAVITY (1.97-2.4 µm) using 1 000 live points. We also explored the additional parameters provided by ForMoSA: the radius (R), radial velocity (rv), and the projected rotational velocity (v.sin(i)). The priors and their associated posteriors are presented in Tables 3 and 4 for the different models. Figure 3 shows the best fit for the four family of atmospheric models. Posterior distributions can be found in the Appendices.
@@
All models faithfully reproduce the observations, accurately capturing the shape of the pseudo-continuum, the pattern of H2O and CO absorptions, and the absolute flux level. Residuals are relatively low () except at the edges of the wavelength interval (2.02m; 2.35m) where telluric residuals are more prominent (see Fig. 14). Comparing the Bayesian evidences (see Table 5 below), Exo-REM and Sonora are statistically preferred among the four models (see Sect. 3 for more details). The Bayesian evidence for the ATMO model fit is significantly lower. This model proved to be the least able to give physical constraints on the planet’s physical parameters, with parameters converging at the border of the grid.
@@ 4.1 Physical and atmospheric parameters
For the effective temperatures, except ATMO, all models give consistent values between and K. These values are slightly lower than the ones obtained by GRAVITY Collaboration et al. (2020) using a lower resolution K-band spectrum ( K for Exo-REM and K for petitRADTRANS444https://petitradtrans.readthedocs.io/en/latest/ Mollière et al., 2019). They are also lower than those predicted by evolutionary models (Chilcote et al., 2017, K).
For surface gravity, evolutionary models predict values for Pic b of dex (Chilcote et al., 2017). Exo-REM and Sonora retrieve larger values of dex and dex, respectively, while BT-Settl is noticeably lower at dex. ATMO also converges at the border of the considered grids, meaning that the model likely has difficulties to properly represent the pseudo-continuum at K-band.
@@
The carbon-to-oxygen ratio obtained with Exo-REM is solar () and super-solar for BT-Settl (). These values differ significantly from those adopted by GRAVITY Collaboration et al. (2020) (), Landman et al. (2024) (), and Worthen et al. (2024) (), but is in agreement with the recent result of Kiefer et al. (2024) () and Houllé et al. (2024, subm.) (), both using Exo-REM.
The cloud sedimentation efficiency parameter fsed is explored only by the Sonora model grids. This parameter ranges from 1 (fully cloudy atmosphere) to 8 (no clouds). In our analysis, we obtain an upper constraint of f, indicating a cloud-rich atmosphere for this object, probably form by silicate and/or metal-oxide condensates as expected from L/T objects surveys (Petrus et al. in prep). This result is somewhat surprising, as Sonora predicts high metallicity, with metals typically trapped in clouds for such objects. Additionally, the adiabatic index , explored only by ATMO, converged towards the upper edge of the grid with . A very high adiabatic index suggests a strong temperature gradient in the atmosphere, which could simulate the presence of clouds. This finding is consistent with the low sedimentation efficiency obtained using Sonora.

### v2

@@ 2.2 Archival data
In addition to GRAVITY data, we used archival spectroscopic and photometric data. Table 2 offers more details on the characteristics of each observations. Data reduction and post-processing analysis can be found in the referenced paper of each study. The specific analysis of high-resolution data is presented in more details in Sect. 6. We did not include the six new JWST/NIRCam photometric points presented in Kammerer et al. (2024), as the two PSF subtraction methods used in that work produced significant scaling differences in the final flux values. Moreover, we did not expect these additional points to contribute meaningfully to the fit in our current weighted scheme.
@@
| VLT/SPHERE | YJ-band spectroscopy | 0.95–1.35 µm | 50 | 22.5 | yes | (*) |
@@
| Gemini/GPI | YJHK-band spectroscopy | 0.9–2.4 µm | 45 | 10 | yes | (b) |
| VLTI/GRAVITY | K-band spectroscopy | 1.97–2.4 µm | 4,000 | 20 | yes | (**) |
| VLT/SINFONI | K-band spectroscopy | 1.929–2.472 µm | 4,120 | 20 | no | (c) |
| VLT/CRIRES+ | K-band spectroscopy | 2.06–2.47 µm | 100,000 | ¿5 | no | (d, e) |
| M-band spectroscopy | 3.51–5.22 µm | 100,000 | ¿5 | no | ||
| VLTI/MATISSE | LM-band spectroscopy | 2.76–4.99 µm | 387–700 | 35 | yes | (f) |
| JWST/MIRI | M-band spectroscopy | 4.90–6.8 µm | 30–40 | 15 | yes | (g) |
We divided our analysis in tree main parts. In Sect. 4, we focus on the analysis of the new GRAVITY spectra. In Sects. 5 and 6, we describe the combination of this spectrum with the available datasets presented in Sect. 2. All of the observational data, together with the complete set of models and results is available on Zenodo555https://zenodo.org/records/17609314.
@@ 3 Methods
In this section, we present our fitting strategy based on forward modeling, as well as our Bayesian ForMoSA code. Sects. 3.1 and 3.2, respectively, code with the forward models we used and the new additions made to ForMoSA for the purposes of this work.
To extract information from exoplanet spectra, a common approach is to compare them with synthetic spectra generated from models of atmosphere. The comparison heavily relies on Bayesian inversion methods that can be divided into two categories: forward modeling and retrieval. The forward modeling approach (Samland et al., 2017) works ”from the model to the data” by using a pre-computed grid of several thousand of spectra and comparing them to the data using Bayesian inference methods. This approach has the advantage of being ”self-consistent,” meaning that it faithfully maintains the physical interpretation of the observed spectrum, as each synthetic spectrum in the grid is computed with strong physical priors. It is also relatively fast since you do not need to generate a new model at each step, but only need to re-interpolate the grid. The main drawbacks are the dependencies on both the quality and the size of the parameter space proposed by the model grid, as well as the systematic deviations from the models that can arise from a lack of physics or hidden fixed free-parameters in the forward model.
Each model is compared iteratively using Bayesian inversion techniques. The goal is to estimate the posterior probability p() of the model parameters given the data and prior knowledge on the parameters pd, using Bayes theorem,
@@
p is the probability of the data given a parameter set, also known as the likelihood function. This is the function we want to maximize during the inversion.
p is the Bayesian evidence, computed by marginalizing over the parameter space. It can be written as ddX, with X as the prior volume.
The Bayesian evidence is especially useful when comparing two competing models, and . Unlike the likelihood , it is independent of any single set of parameters as it integrates over all parameters (and is therefore more general in that sense). Concretely, if we label and as the Bayesian evidence of the two competing models, we can compute the Bayes factor using
Typically, a Bayes factor above 12 (or above 2.5) can be interpreted as a ”moderate need” for the first model to describe the data instead of the second ( significance, see Table 4 of Benneke & Seager, 2013). detection. Recently, Kipping & Benneke (2025) demonstrated that this commonly used correspondence between and can lead to overestimations of the actual detection significance. For completeness, we report both metrics, but we rely primarily on the Bayesian evidence as the main diagnostic tool for the model comparison. This approach will be used in Sects. 4.3 and 5 to assess the significance of the 13CO detection. For a more detailed description of the Bayesian inference methods, we refer to Knuth et al. (2015).
For this work, we chose to use the forward modeling approach with our Bayesian inversion tool, ForMoSA. A detailed description of the code can be found in Petrus et al. (2020), Petrus et al. (2021), Petrus et al. (2023), and Petrus et al. (2024). In brief, ForMoSA uses a grid of forward models (see Sect. 3.1) along with spectroscopic and/or photometric data. The tool then proceeds to adapt these models to the data (in terms of wavelength range coverage, resolution, continuum, etc.) to make them comparable. The inversion is performed by maximizing the likelihood function using the nested sampling method (see Skilling, 2004). The grid is interpolated linearly in the N dimension of the parameter space using the python packaged xarray666https://docs.xarray.dev/en/stable/.
@@ 3.1 Forward models
Atmospheric models aim to replicate the emission spectra of exoplanets. To do this, they typically compute a blackbody-like emission spectrum from an interior model, which is then modified by a radiative-convective atmosphere composed of various opacity sources (and sometimes clouds), following a parameterized pressure-temperature profile; all within a simplified 1D framework. These models cover different wavelength ranges and spectral resolutions, as illustrated in Fig. 1. We give a brief summary of the grids we used below, previously presented in Petrus et al. (2024).
@@
Exo-REM (Charnay et al., 2018) is our reference model. It includes non-equilibrium chemistry of CO–CH4 and N2–NH3 and takes into account multiple opacity sources: H2–H2 and H2–He collisions induced absorptions, H2O, CH4, CO, CO2, NH3, PH3, TiO, VO, and FeH ro-vibrationnal transitions as well as Na and K resonant lines. Clouds formation based on Si, Fe, Na2S, KCl, and H2O species was considered, but with a simplified cloud micro-physics (i.e., only the fastest chemical processes were taken into account). We used a custom version of the grid to explore the abundance of the isotope 13C resulting in these five free parameters: [1200 K; 2000 K], log(g) [3.5 dex; 4.5 dex], [M/H] [-0.5 dex; 1 dex], C/O [0.40; 0.65], and log(12C/13C) [1.0; 2.6]. Exo-REM is known to produce deviant spectra so we performed a simple cleanup by computing the integrated fluxes across the grids and removing the spectra that did not follow the approximate Stefan–Boltzmann law. Furthermore, we chose to remove all nodes at = 1500 K from the grid, as they appeared to have some deviant spectra driving the solutions.
@@
ATMO (Tremblin et al., 2015) are a family of cloudless models. They explore numerous opacity sources : H2–H2 and H2–He collision induced absorptions, NH3, H2O, CH4, CO, TiO, and VO ro-vibrationnal transitions as well as Na and K resonant lines. These models include two non-equilibrium chemical reactions: CO–CH4 and N2–NH3 that will drive fingering convection in the atmosphere and change the temperature gradient. This will uniquely emulate the reddening effect of clouds. We also utilized a custom version of this grid exploring 13C. The parameter space we explored is: [1600 K; 1900 K], log(g) [3.5 dex; 4.5 dex], [M/H] [-0.6; 0.6], [1.01; 1.05], and log(12C/13C) [1.0; 2.6] ( is the adiabatic index driving the temperature gradient). The grids were set with C/O = 0.55. In Sect. 4, we use an extended grid in and log(g) that explored the following parameters : [800 K; 3000 K], log(g) [2.5 dex; 4.5 dex], [M/H] [-0.6; 0.6], and C/O [0.3; 0.7].
@@
BT-Settl (Allard et al., 2012) model grids explore more complex cloud microphysics than Exo-REM and include non-equilibrium chemistry (for CO, CH4, N2, NH3, and CO2) as well as vertical mixing. To properly take into account the cloud micro-physics, BT-Settl divides the atmosphere into multiple layers and computes the distribution and size of grains by comparing their characteristic times of condensation, coalescence, dispersion, and sedimentation in the atmosphere. These clouds are Fe and Si based. We used a custom version of this grid exploring the C/O, giving the following parameter space: [1400 K; 2200 K], log(g) [2.5; 5.0], and C/O [0.2754; 1.096]. The metallicity is solar () for this grid.
@@
Sonora (Morley et al., 2024) models include vertical mixing, multiple opacity sources and parameterized clouds. We used the custom Diamondback sub-model grid that contains, in total, 31 molecules, atoms, and ions as opacity sources (hydrocarbons, CO, CO2, OCS, Li, K, Na, LiH, LiF), as well as multiple collision induced absorptions (H2–H2, H2–He, H2–N2, H2–CH4). Clouds were parameterized using Ackerman & Marley (2001) approach, meaning that their opacities were computed for Mie diffusion on MgSiO3, Mg2SiO4, and Al2O3. In the end, the model grid effectively explores the following free parameters: [900 K; 2400 K], log(g) [3.5 dex; 5.5 dex], [M/H] [-0.5; 0.5], and f [1; 8] (f is the sedimentation efficiency parameter). The grids were set with C/O = 0.55.
@@
Atmospheric retrievals and forward modeling approaches both deal with a number of issues when combining different observations into a single Bayesian framework. A non-exhaustive list of theses limitations include: data that spectraly overlap (i.e., GPI and GRAVITY); data that have different noise estimates (i.e., spectral covariances or not); data that are not calibrated in flux or have inaccurate flux calibration; data that are taken at different epochs and spectra for which the spectral continuum can not be trusted or has been removed at previous analysis steps (Doppler spectroscopy, etc). Inspired by the work of Hayoz et al. (2023), we incorporated a Multimodal Option for Simultaneous Analysis and Improved Constrains (MOSAIC) in ForMoSA aim to reduce theses limitation by adapting and inverting the different datasets in parallel.
The general idea is summarized on Fig. 2. Each dataset was first imported and adapted alongside a grid of synthetic spectra taken from a single atmospheric model. This means that the continuum subtraction and/or wavelength windows can be applied separately for each data-grid pair and then stored for the inversion. Secondly, each dataset was compared with the interpolated grid using separate likelihood functions during the inversion. The grids were modified by both interpolating them with the original model parameters and then adding transformations with extra-grid parameters, , that can vary from dataset to dataset (i.e., different scaling factors, ck, radial velocities, rv, projected rotational velocities, , etc.). In this way, we were able to control the transformations we want to apply on a given data+grid pair. The stored sub-grids are usually much smaller than the original grids. Therefore, this step is relatively fast.
New likelihood functions are now available to account for differences in the data reduction scheme and, in particular, noise processing. Given a set of data (spectrum or photometry), , its associated error bars, or covariance matrix, C, and the model vector, m, the classical distribution can be defined as
where the observational data points are independent. In addition, we implemented
when the data covariances are provided as input of the code. We also allowed for various proposed mappings (Zucker, 2003; Brogi & Line, 2019; Ruffio et al., 2019) of cross-correlation signals to likelihood (CCF-to-log), which can be used when Doppler spectroscopy (e.g., Keck/KPIC, VLT/HiRISE, etc) was performed solely or in addition to regular spectrophotometric observations (e.g., JWST, VLT/SPHERE, Gemini/GPI, etc.).
We computed the joint likelihood function , summing over the likelihoods corresponding to each dataset taken as input. This means that this upgraded version of ForMoSA does not account for correlations either between datasets from the same instrument taken at different epochs or between datasets from different instruments. As is common practice in the field, each dataset was treated as statistically independent.
@@ 4 Inversion of GRAVITY spectra
For this inversion, the likelihood function was computed using the generalized distribution (see Sect. 3.2) taking as input the covariance matrixes provided by the GRAVITY pipeline. The inversion was performed across the entire spectral coverage of GRAVITY (1.97–2.4 µm) using 1 000 live points. We also explored the additional parameters provided by ForMoSA: the radius (R), radial velocity (rv), and the projected rotational velocity (). The priors and their associated posteriors are presented in Tables 7 and 8 for the different models. Figure 3 shows the best fit for the four family of atmospheric models. Posterior distributions can be found in the Appendices.
@@
All models faithfully reproduce the observations, accurately capturing the shape of the pseudo-continuum, the pattern of H2O and CO absorptions, and the absolute flux level. Residuals are relatively low () except at the edges of the wavelength interval (m; m) where telluric residuals are more prominent (see Fig. 14). Comparing the Bayesian evidence (see Table 9 below), Exo-REM and Sonora are statistically preferred among the four models (see Sect. 3 for more details). The Bayesian evidence for the ATMO model fit is significantly lower. This model proved to be the least able to give physical constraints on the planet’s physical parameters, with parameters converging at the border of the grid.
@@ 4.1 Physical and atmospheric parameters
For the effective temperatures, except ATMO, all models give consistent values between and K. These values are slightly lower than the ones obtained by GRAVITY Collaboration et al. (2020) using a lower resolution K-band spectrum ( K for Exo-REM and K for petitRADTRANS101010https://petitradtrans.readthedocs.io/en/latest/ Mollière et al., 2019). They are also lower than those predicted by evolutionary models (Chilcote et al., 2017, K).
For surface gravity, evolutionary models predict values for Pic b of dex (Chilcote et al., 2017). Exo-REM and Sonora retrieve larger values of dex and dex, respectively, while BT-Settl is noticeably lower at dex. ATMO also converges at the border of the considered grids, meaning that the model likely has difficulties to properly represent the pseudocontinuum in the K band.
@@
The carbon-to-oxygen ratio obtained with Exo-REM is solar () and super-solar for BT-Settl (). These values differ significantly from those adopted by GRAVITY Collaboration et al. (2020) (), Landman et al. (2024) (), and Worthen et al. (2024) (), but is in agreement with the recent result of Kiefer et al. (2024) () and Houllé et al. (2025) (), both using Exo-REM.
The cloud sedimentation efficiency parameter, f, is explored only by the Sonora model grids. This parameter ranges from 1 (fully cloudy atmosphere) to 8 (no clouds). In our analysis, we obtain an upper constraint of f, indicating a cloud-rich atmosphere for this object, probably form by silicate and/or metal-oxide condensates as expected from L/T objects surveys (Petrus et al., 2025). This result is somewhat surprising, as Sonora predicts high metallicity, with metals typically trapped in clouds for such objects. Additionally, the adiabatic index , explored only by ATMO, converged towards the upper edge of the grid with . A very high adiabatic index suggests a strong temperature gradient in the atmosphere, which could simulate the presence of clouds. This finding is consistent with the low sedimentation efficiency obtained using Sonora.